40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9606     Entry time: Wed Feb 5 20:41:57 2014
Author: Den 
Type: Update 
Category: LSC 
Subject: calibrated spetra from OAF test 

We did online adaptive filtering test with IMC and arms 1 year ago (log 7771). In the 40m presentations I can still see the plot with uncalibrated control spectra that was attached to that log. Now it the time to attach the calibrated one.

Template is in the /users/den/oaf

Attachment 1: oaf_cal.pdf  94 kB  | Show | Show all
ELOG V3.1.3-